OpenStudy (anonymous):

The product of two consecutive positive even integers is 12,320. What are the two integers?

OpenStudy (anonymous):

We can use an equation where the first number is x and the second number is x+2 so we have x(x+2) = 12320 then multilply on the left-side x^2 + 2x = 12320 and subtract 12320 on both sides x^2 + 2x - 12320 = 0 Do you know the next step?

OpenStudy (anonymous):

divide both sides by 2?

OpenStudy (anonymous):

Not quite ... factor ..... x^2 + 2x - 12320 = (x + 112)(x - 110) So we have (x + 112)(x - 110) = 0 Do you know the next step?

OpenStudy (anonymous):

combine like terms?

OpenStudy (anonymous):

solve each factor x + 112 = 0 or x = -112 but this is not positive x - 110 = 0 or x =110 The answers are 110 and 112

OpenStudy (anonymous):

Great! thank you so much for explaining your steps!